OpenStudy (baldymcgee6):

Spike (who is 6 feet tall) stands at a distance of x feet from a 24 foot high street light. Express the length L of his shadow as a function of x.

OpenStudy (baldymcgee6):



OpenStudy (baldymcgee6):

@Hero, @hartnn

OpenStudy (baldymcgee6):

similar triangles?

hartnn (hartnn):

know the concept of similar triangles?

hartnn (hartnn):

yes,they are similar,can u prove it?

OpenStudy (baldymcgee6):

they share the same hypontenuse.

hartnn (hartnn):

nopes,they don't.

OpenStudy (baldymcgee6):

then i dont know

hartnn (hartnn):

try proving two pairs of angles as equal,one is obvious.

OpenStudy (baldymcgee6):

right, they share the same right angle and the same acute angle.

OpenStudy (baldymcgee6):

24/(x+L)= 6/L

hartnn (hartnn):

u got everything! good work. now just isolate L.

OpenStudy (baldymcgee6):

how do i isolate L?

hartnn (hartnn):

first cross multiply.

OpenStudy (baldymcgee6):

24L=6x+6L

hartnn (hartnn):

or write 24=6*4 and cancel 6

OpenStudy (baldymcgee6):

L=6x/18

OpenStudy (baldymcgee6):

L=x/3

hartnn (hartnn):

yup,correct :)